How many liters of a 50% acid solution must be mixed with a 20% acid solution to get 240 L of a 40% acid solution.
How many liters of a 50% acid solution must be used?

Found 2 solutions by ewatrrr, richwmiller:
Answer by ewatrrr(24785)

.50x + .20(240L-x) = .40(240L)
x = .20(240L)/.30, liters of a 50% acid solution

Answer by richwmiller(17219)

a+b=240,
0.5*a+0.2*b=0.4*240
a=240-b
0.5*(240-b)+0.2*b=96
120-0.5b+0.2*b=96
-0.3*b=-24
b=80
a=240-b
a=160 at 50% b=80 at 20%
check
0.5*160+0.2*80=0.4*240
80+16=96
96=96
ok